OpenStudy (anonymous):

PLEASE HELP MEEE!!! True or False. 0 has an odd multiplicity the polynomial x^4-25x^2

OpenStudy (anonymous):

anyone???

jimthompson5910 (jim_thompson5910):

I think it's asking if x = 0 is a root with odd multiplicity

jimthompson5910 (jim_thompson5910):

If so, then notice how x^4-25x^2 = 0 can factor to x^2(x^2-25) = 0

jimthompson5910 (jim_thompson5910):

Can you see what the roots are?

OpenStudy (anonymous):

x=0 x=+5 x=-5??

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

x^2 = 0 yields two roots that are the same, which are x = 0 So the multiplicity of x = 0 is two

OpenStudy (anonymous):

so it is even which would make it false??

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok thanks can you help me with a few others??

jimthompson5910 (jim_thompson5910):

yw

jimthompson5910 (jim_thompson5910):

post each question one at a time to the left

jimthompson5910 (jim_thompson5910):

and let me know you've done so if I don't come to it